OpenStudy (anonymous):

simplify (cos4x-cos2x)/(sin4x+sin2x) the answer choices are : cotx, tanx, cotxtanx, or 1

OpenStudy (anonymous):

i believe it is none of the above.

OpenStudy (anonymous):

think it is -tangent, let me check with pencil and paper

OpenStudy (anonymous):

That's what Wolfram says...

OpenStudy (anonymous):

The top is -Sin^2 2x so if we can get the cos on the bottom, we're done.

OpenStudy (anonymous):

damn this must be complicated algebra day. you are going to have to rewrite all this in terms of sin(x) and cos(x)

OpenStudy (anonymous):

Yeah, the bottom is a bit of a pain...

OpenStudy (anonymous):

@estudier top is \[\cos^2(2x)-\sin^2(2x)-\cos(2x)\] and i guess you have to reduce again yes?

OpenStudy (anonymous):

Yes but Cos^2 2x is Co2 x -Sin^2x

OpenStudy (anonymous):

i need more scrap paper

OpenStudy (anonymous):

You really ought to be able to do something with the symmetry here...

OpenStudy (anonymous):

I must be doin' something daft...

OpenStudy (anonymous):

ok i give up. to much damned algebra for me.

OpenStudy (anonymous):

I'll get there in the end...

OpenStudy (anonymous):

cheat, graph the thing and see what you get

OpenStudy (anonymous):

I'm pretty sure it's tan, it feels like tan, not so sure about the sign, though..

OpenStudy (anonymous):

Think I will have to switch sin for -cos

OpenStudy (anonymous):

Was finally able to find a common factor to reduce the sob. Wolfram is right, it's -tanx.(it's definitely not 1)